Find two numbers whose sum is -6 and whose difference is -2 (please show steps)
Answer by mananth(16946)

let one number be x
other number be y
x+y=-6
x-y=-2
Add the two equations to eliminate y
(x+y)+(x-y) = (-6)+(-2)
x+y+x-y = -8
2x= -8
x= -8 /2
x= -4
x-y = -2
-4-y=-2
-y=4-2
-y= 2
v= 2/-1
y= -2